OpenStudy (anonymous):

Pablo is making ramps of different lengths. The ramps are made with rectangular and square wooden boards. The relationship between the length of the ramp in feet, f(n), and the number of square boards used, n, is represented by the function below. f(n) = 20n + 3 , where 2 ≤ n ≤ 6 Which statement is correct about the range of the length of the ramp for the given domain?

OpenStudy (anonymous):

ell the range is the numerical vaules for y that fit within our domain (x) of our function. So when we plug in 2 to f(n), we get 43 and when we plug in 6 to f(n) we get 123.

OpenStudy (anonymous):

Well, you have to sustitute the values of n for n = 2 f(n) = 20(2) + 3 f(n) = 43 for n = 6 f(n) = 20(6) + 3 f(n) = 123 So the answers is A.It varies from 43 to 123.

OpenStudy (anonymous):

2(20) + 3 = 43 is the lowest range possible. 6(20)+3 = 123 is the highest range possible. A) it varies from 43 to 123.